# On Improvisation and Open-Endedness: Insights for Experiential AI

### v1

@@ Introduction
Improvisation is a foundational mode of human creativity, evident across diverse art forms—from jazz music to contemporary dance to everyday conversation. It involves the spontaneous generation and adaptation of action in real-time, without a predetermined script. Improvisational arts have a long-standing presence in human culture: in music, improvisation has been widely studied—from instrumental jazz solos to freestyle rap battles (Berliner 2010; Torrance and Schumann 2019; Lewis and Piekut 2016); in dance, forms like contact improvisation—originated by Steve Paxton in 1972 (1975)—emphasizes movement through physical contact; even emerging digital practices such as live coding treat code as a real-time performative medium (Selvaraj et al. 2021). Though it does not always produce brilliant works of art, improvisation is both experiential and open-ended by nature—it responds to the present moment, performing unrepeatable actions and generating continuous novelty—qualities highly valued in artistic creativity (Karlen-Gelli 2024).
@@
In the field of artificial life (ALife), open-endedness (OE) refers to a system’s capacity for ongoing innovation, producing unbounded novelty or endless “interestingness” without converging to a static equilibrium (Packard et al. 2019). In nature, life itself exemplifies an open-ended evolutionary process: over billions of years, biological evolution has continually generated novel forms and behaviors without a fixed endpoint (Taylor 2019). Cultural evolution provides a second real-world example of unbounded innovation (Borg et al. 2024). In contrast, artificial systems—such as digital evolution system or generative AI—often plateauing after an initial burst of novelty (Stanley 2019). Indeed, replicating the continually creative dynamics of natural or cultural evolution has been deemed “the the last grand challenge” (O. Stanley, Lehman, and Soros 2017). Soros et al. (2024) highlights a close relationship between creativity and open-endedness, suggesting that the difficulty of producing genuinely creative systems parallels the difficulty of achieving OE itself. Yet, in the field of computational creativity (CC)—which Colton Simon and Wiggins Geraint A. (2012) argue is “the final frontier”—most machine-generated creations remain closed-ended: produced via pre-programmed rules, fixed objectives, or models trained on past human data. Although recent generative AI models (large language models, video generation models, etc.) can produce impressively creative content and may appear open-ended, they do not truly achieve open-endedness. They often repeat patterns or clichés, essentially regurgitating averages of their training data. The fundamental reason may be a lack of agency: human creativity is not purely objective-oriented—it involves conviction, serendipity, curation, and enjoyment. As Stanley and Lehman (2015) argue in “Why Greatness Cannot Be Planned,” current AI systems lack this self-directed pursuit of endless “interestingness” to explore the unknown or push their creativity beyond their training data and preset objectives. They lack the experiential learning that fuels human creators’ continual growth. As Silver and Sutton (2025) argues, AI development is entering an “era of experience”—moving beyond the “era of human data.” This shift requires agents to learn through real-world interaction and self-directed exploration, rather than simply consuming static datasets. Hughes et al. (2024) even posit that open-endedness is essential for any future artificial superintelligence—such an AI would need a self-driven capacity to discover novelty beyond human knowledge.
@@
Recent developments in human-computer interaction (HCI) and AI research have started exploring improvisational systems, from robots that jam with musicians to virtual partners that dance with humans. For instance, Pataranutaporn et al. (2024) present Human–AI co-dancing performances where virtual dancers with generative models collaborate with live performers, evolving choreographic ideas together. Trajkova et al. (2024)’s LuminAI system allows people to improvise dance movements with an AI dancer that observes, mirrors, and riffs on the user’s gestures. In improvised theatre, Mathewson and Mirowski (2018)’s Improbotics experiment placed a chatbot into a live improv comedy show, with audiences and even fellow actors unaware of which performer was AI. Researchers have also crafted improvisational experiences with non-humanoid agents: Dancing with Drones by Dong et al. (2024), put autonomous drones on stage with human dancers, finding that an “intercorporeal understanding” had to be developed between the choreographer and the drones’ behaviors to co-create expressive movement. Likewise, Drone Chi by La Delfa et al. (2020) explored a Tai Chi-inspired human–drone interaction in which a flying robot and a person respond to each other’s motions in a close, somatic loop. In music, Benford et al. (2024) built an AI folk fiddler and found that performers had to negotiate autonomy and trust – e.g. by pre-setting the AI’s “looseness of control” and using simple intensity cues during performance – so that the human could enter a flow state with the AI and still feel safe to experiment. These pioneering systems hint at the possibilities of Experiential AI that can improvise with humans or other AIs. However, they also underscore the challenges: maintaining “interestingness”, preventing the AI from generating dull or chaotic outputs, and enabling the AI to meaningfully evolve its behavior over time. In short, today’s AI improvisers can simulate improvisation, but to achieve the real qualities of human improvisation, new principles are needed. We see an opportunity to connect the wisdom of improvisational art practice with future open-ended experiential AI in computational creativity. A fascinating question arises:
@@
Improvisation has deep roots in human creative practice. Historically, improvisation in music predates written composition – for example, Western classical figures like Mozart were renowned for improvising at the keyboard, and non-Western traditions (Indian raga, etc.) have rich improvisational frameworks. In jazz, improvisation is a defining feature: it is often described as “spontaneous creative behavior” requiring novel combinations of learned patterns and real-time invention (Berliner 2010). Jazz ensembles showcase collective improvisation, where musicians co-create by listening and responding to each other, producing a delicate interplay of solo and accompaniment. Freestyle rap similarly involves instant composition of lyrics and rhythm; neuroscientific studies highlight that freestyle rap is a “multidimensional form of creativity at the interface of music and language,” engaging brain networks for spontaneous, flow-state performance (Liu et al. 2012). In dance, improvisation ranges from solo movement exploration to collaborative forms like Contact Improvisation, an improvised partner dance initiated by Steve Paxton in 1972. Contact Improv is described as “an open-ended exploration of the kinesthetic possibilities of bodies in contact”, involving shared physical dialogue and real-time co-adaptation of movement (Paxton 1975). Across these domains, human improvisers cultivate skills of attentive listening, quick adaptation, and creative risk-taking. Improvisation is often likened to everyday life: “Life is improvisation”, as comedian Tina Fey quipped, capturing how we navigate daily interactions without a script. Indeed, anthropologist Mary Catherine Bateson noted that “ambiguity is the warp of life…learning to savor the vertigo of doing without answers” allows one to play with patterns and find coherence amid uncertainty (Bateson 1995). In essence, improvisation is a fundamental mode of human creativity – a way to make something new in the present moment by drawing on one’s repertoire, environment, and imagination.
@@
Open-endedness (OE) has become an important concept in ALife and AI research, referring to an open-ended generative process that can produce unbounded novelty over time (Packard et al. 2019). In ALife, open-ended evolution (OEE) was identified over two decades ago as a grand challenge: “to determine what is inevitable in the open-ended evolution of life” (Soros et al. 2024). Natural evolution never settles in a static state – it continually produces new species, forms, and behaviors. Researchers have attempted to recreate such dynamics in virtual worlds (e.g. Tierra, Avida), but so far even the most complex simulations eventually run out of novel variations and reach a limit. Bedau et al. (2000) noted that achieving indefinite evolutionary innovation in artificial systems is inevitable yet elusive. Definitions of OEE vary: for instance, Taylor (Taylor 2019) defined it as a system that keeps evolving new forms instead of converging to an optimum, while others emphasize an “indefinite increase in complexity” or continual adaptive novelty. O. Stanley, Lehman, and Soros (2017) argued that solving open-endedness is “the last grand challenge” for AI, on par with – and perhaps integral to – achieving human-level AI. In parallel, creativity in AI as the goal of Computational Creativity (CC) has been called AI’s “final frontier” (Colton Simon and Wiggins Geraint A. 2012). Both creativity and OE are central, enduring research themes, yet neither is close to solved. Indeed, the two are deeply linked: creativity is a hallmark of human and animal cognition, while open-ended innovation is foundational to life’s evolution. Researchers have proposed that truly creative AI may require open-ended generative processes rather than fixed problem-solving (Soros et al. 2024). Efforts to formalize “interestingness” and intrinsic novelty have appeared in both fields – e.g. Schmidhuber (1991)’s theory of curiosity-driven learning (rewarding novelty or surprise), Pathak et al. (2017)’s curiosity modules in reinforcement learning, or Secretan et al. (2008)’s interactive evolution of images – but such metrics only capture narrow aspects of what humans find creative. Purely quantitative measures of complexity or interestingness often fail to match the nuanced human sense of what is truly creative. The challenge, then, is not only to generate novelty, but to ensure an interestingness that is sustainable and meaningful. Recent work argues that open-endedness might be essential for advanced AI: Hughes et al. (2024) posit that any artificial super-intelligence must continually self-improve and discover novelty to surpass human abilities. In practical terms, this means moving beyond training on static human datasets. As Silver and Sutton (2025) observe, high-quality human data is finite and nearly exhausted for current AI systems – progress from human examples is slowing, necessitating a new paradigm. Open-ended systems address this by enabling agents to generate their own experience and thereby their own ever-growing training data. In summary, open-endedness and interestingness form a guiding vision for next-generation AI: systems that, like living or creative processes, can grow autonomously, continually yielding new, interesting outcomes without end.
@@ Results
Every expert we interviewed underscored the need to suspend judgment during improvisation in order to enter a flow state. P4 put it succinctly: “When you’re improvising, you can’t be judging every move as ‘good’ or ‘bad’ – if you do, you hesitate and the flow is gone. You have to treat everything as an offer and run with it.” This ethos of radical acceptance is taught in improv theater as the “Yes, and” rule – you accept whatever your partner or the situation provides, and build upon it. In dance and music, it manifests as not second-guessing oneself. P6 noted that even if she makes an awkward move in contact improv, she immediately “recycles” it into the next movement rather than stopping in embarrassment. This creates an unbroken continuity: the improvisation may twist in unexpected directions, but it doesn’t halt. The concept of flow (Nakamura and Csikszentmihalyi 2014) – a state of absorbed focus where action and awareness merge – came up repeatedly. Improvisers described flow experiences where “time dilates and self-consciousness disappears.” Neuroscience corroborates this: during improvisation, brain regions associated with self-monitoring and evaluation deactivate, supporting a state of free expression (Limb and Braun 2008). To achieve this, improvisers deliberately quiet their inner critic. P5 said, “I tell my students to improvise as if nobody’s watching, even if on stage – that freedom is where the real ideas come.” Interestingly, this non-judgment extends to how improvisers treat others in group improv: there’s a norm of not criticizing or blocking another’s contribution, but rather finding a way to make it work. This creates a safe space where creativity can flourish without fear.
Implication
@@
Great improvisations often feel like they are walking a tightrope between order and chaos. Our interviewees described it as riding the edge: “If it’s too predictable, it gets boring; if it’s too random, it falls apart,” said P4. The interestingness sweet spot lies in between – what complexity science calls the edge of chaos, where a system has enough stability to maintain coherence but enough instability to generate novelty . P2 (musician) gave a musical example: in jazz, improvisers sometimes go “outside” the chord changes (deliberately dissonant or offbeat) and then resolve back “inside,” creating tension and release. “It’s like we stretch the rubber band of the music but don’t let it snap,” he explained. Similarly, P1 (dancer) mentioned playing with the boundaries of control: moving just beyond her comfort zone (faster, or with a new quality) so that she has to adapt creatively, but not so far that she loses all technique. Interestingly, one contact improviser (P4) articulated two personal principles: self-sustain and irreversibility. By self-sustain, he meant each dancer should carry their own weight and balance (a grounded structure) so they are not overly dependent; by irreversibility, he meant that once a movement is done, they don’t try to undo it or rewind – they fully commit and move to the next. “All your actions should be irreversible – that gives the maximum potential for the next move,” he said, tying it to the idea that if you never try to simply cancel what happened, you keep the forward momentum at the edge of chaos. Improvisers also often impose light constraints or structures on themselves to shape the chaos. P3 (live coder) might decide to restrict to a certain scale or algorithm during a segment, using that as a launchpad for creativity. These self-imposed rules create a framework that prevents total randomness while still demanding invention within new bounds (a concept analogous to exploring adjacent possible in creative theory).
Implication
@@ Conclusion
Improvisation in human artistry offers a rich tapestry of insights for crafting the next generation of interactive, experiential AI systems. These human-centric qualities can inform concrete design principles for AI. In essence, we envision Experiential AI improvisers that are not just content generators but partners – with a degree of presence and creativity that makes interacting with them feel alive, interesting, and continuously meaningful. By grounding future AI improvisers in the hard-won wisdom of ancient improvisational practices and the spirit of life’s improvisational nature, we can move closer to AI that doesn’t merely output on command, but improvises – co-creating endless interesting possibilities hand-in-hand with humans.

### v2

@@ Introduction
Improvisation is a foundational mode of human creativity, evident across diverse art forms—from jazz music to contemporary dance to everyday conversation. It involves the spontaneous generation and adaptation of action in real-time, without a predetermined script. Improvisational arts have a long-standing presence in human culture: in music, improvisation has been widely studied—from instrumental jazz solos to freestyle rap battles (Berliner2010Thinking; Torrance2019spur; Lewis2016Oxford); in dance, forms like contact improvisation—originated by Steve Paxton in 1972 (Paxton1975Contacta)—emphasizes movement through physical contact; even emerging digital practices such as live coding treat code as a real-time performative medium (Selvaraj2021Live). Though it does not always produce brilliant works of art, improvisation is both experiential and open-ended by nature—it responds to the present moment, performing unrepeatable actions and generating continuous novelty—qualities highly valued in artistic creativity (KarlenGelli2024Composing).
@@
In the field of artificial life (ALife), open-endedness (OE) refers to a system’s capacity for ongoing innovation, producing unbounded novelty or endless “interestingness” without converging to a static equilibrium (Packard2019OpenEnded). In nature, life itself exemplifies an open-ended evolutionary process: over billions of years, biological evolution has continually generated novel forms and behaviors without a fixed endpoint (Taylor2019Evolutionary). Cultural evolution provides a second real-world example of unbounded innovation (Borg2024Evolved). In contrast, artificial systems—such as digital evolution system or generative AI—often plateauing after an initial burst of novelty (Stanley2019Why). Indeed, replicating the continually creative dynamics of natural or cultural evolution has been deemed “the the last grand challenge” (stanley2017openendedness). Soros2024Creativityb highlights a close relationship between creativity and open-endedness, suggesting that the difficulty of producing genuinely creative systems parallels the difficulty of achieving OE itself. Yet, in the field of computational creativity (CC)—which ColtonSimon2012Computational argue is “the final frontier”—most machine-generated creations remain closed-ended: produced via pre-programmed rules, fixed objectives, or models trained on past human data. Although recent generative AI models (large language models, video generation models, etc.) can produce impressively creative content and may appear open-ended, they do not truly achieve open-endedness. They often repeat patterns or clichés, essentially regurgitating averages of their training data. The fundamental reason may be a lack of agency: human creativity is not purely objective-oriented—it involves conviction, serendipity, curation, and enjoyment. As Stanley2015Whya argue in “Why Greatness Cannot Be Planned,” current AI systems lack this self-directed pursuit of endless “interestingness” to explore the unknown or push their creativity beyond their training data and preset objectives. They lack the experiential learning that fuels human creators’ continual growth. As silver2025welcome argues, AI development is entering an “era of experience”—moving beyond the “era of human data.” This shift requires agents to learn through real-world interaction and self-directed exploration, rather than simply consuming static datasets. Hughes2024Position even posit that open-endedness is essential for any future artificial superintelligence—such an AI would need a self-driven capacity to discover novelty beyond human knowledge.
@@
Recent developments in human-computer interaction (HCI) and AI research have started exploring improvisational systems, from robots that jam with musicians to virtual partners that dance with humans. For instance, Pataranutaporn2024HumanAI present Human–AI co-dancing performances where virtual dancers with generative models collaborate with live performers, evolving choreographic ideas together. Trajkova2024Exploring’s LuminAI system allows people to improvise dance movements with an AI dancer that observes, mirrors, and riffs on the user’s gestures. In improvised theatre, Mathewson2018Improbotics’s Improbotics experiment placed a chatbot into a live improv comedy show, with audiences and even fellow actors unaware of which performer was AI. Researchers have also crafted improvisational experiences with non-humanoid agents: Dancing with Drones by Dong2024Dances, put autonomous drones on stage with human dancers, finding that an “intercorporeal understanding” had to be developed between the choreographer and the drones’ behaviors to co-create expressive movement. Likewise, Drone Chi by LaDelfa2020Drone explored a Tai Chi-inspired human–drone interaction in which a flying robot and a person respond to each other’s motions in a close, somatic loop. In music, Benford2024Negotiating built an AI folk fiddler and found that performers had to negotiate autonomy and trust – e.g. by pre-setting the AI’s “looseness of control” and using simple intensity cues during performance – so that the human could enter a flow state with the AI and still feel safe to experiment. These pioneering systems hint at the possibilities of Experiential AI that can improvise with humans or other AIs. However, they also underscore the challenges: maintaining “interestingness”, preventing the AI from generating dull or chaotic outputs, and enabling the AI to meaningfully evolve its behavior over time. In short, today’s AI improvisers can simulate improvisation, but to achieve the real qualities of human improvisation, new principles are needed. We see an opportunity to connect the wisdom of improvisational art practice with future open-ended experiential AI in computational creativity. A fascinating question arises:
@@
Improvisation has deep roots in human creative practice. Historically, improvisation in music predates written composition – for example, Western classical figures like Mozart were renowned for improvising at the keyboard, and non-Western traditions (Indian raga, etc.) have rich improvisational frameworks. In jazz, improvisation is a defining feature: it is often described as “spontaneous creative behavior” requiring novel combinations of learned patterns and real-time invention (Berliner2010Thinking). Jazz ensembles showcase collective improvisation, where musicians co-create by listening and responding to each other, producing a delicate interplay of solo and accompaniment. Freestyle rap similarly involves instant composition of lyrics and rhythm; neuroscientific studies highlight that freestyle rap is a “multidimensional form of creativity at the interface of music and language,” engaging brain networks for spontaneous, flow-state performance (Liu2012Neural). In dance, improvisation ranges from solo movement exploration to collaborative forms like Contact Improvisation, an improvised partner dance initiated by Steve Paxton in 1972. Contact Improv is described as “an open-ended exploration of the kinesthetic possibilities of bodies in contact”, involving shared physical dialogue and real-time co-adaptation of movement (Paxton1975Contacta). Across these domains, human improvisers cultivate skills of attentive listening, quick adaptation, and creative risk-taking. Improvisation is often likened to everyday life: “Life is improvisation”, as comedian Tina Fey quipped, capturing how we navigate daily interactions without a script. Indeed, anthropologist Mary Catherine Bateson noted that “ambiguity is the warp of life…learning to savor the vertigo of doing without answers” allows one to play with patterns and find coherence amid uncertainty (Bateson1995Peripheral). In essence, improvisation is a fundamental mode of human creativity – a way to make something new in the present moment by drawing on one’s repertoire, environment, and imagination.
@@
Open-endedness (OE) has become an important concept in ALife and AI research, referring to an open-ended generative process that can produce unbounded novelty over time (Packard2019OpenEnded). In ALife, open-ended evolution (OEE) was identified over two decades ago as a grand challenge: “to determine what is inevitable in the open-ended evolution of life” (Soros2024Creativityb). Natural evolution never settles in a static state – it continually produces new species, forms, and behaviors. Researchers have attempted to recreate such dynamics in virtual worlds (e.g. Tierra, Avida), but so far even the most complex simulations eventually run out of novel variations and reach a limit. Bedau2000Open noted that achieving indefinite evolutionary innovation in artificial systems is inevitable yet elusive. Definitions of OEE vary: for instance, Taylor (Taylor2019Evolutionary) defined it as a system that keeps evolving new forms instead of converging to an optimum, while others emphasize an “indefinite increase in complexity” or continual adaptive novelty. stanley2017openendedness argued that solving open-endedness is “the last grand challenge” for AI, on par with – and perhaps integral to – achieving human-level AI. In parallel, creativity in AI as the goal of Computational Creativity (CC) has been called AI’s “final frontier” (ColtonSimon2012Computational). Both creativity and OE are central, enduring research themes, yet neither is close to solved. Indeed, the two are deeply linked: creativity is a hallmark of human and animal cognition, while open-ended innovation is foundational to life’s evolution. Researchers have proposed that truly creative AI may require open-ended generative processes rather than fixed problem-solving (Soros2024Creativityb). Efforts to formalize “interestingness” and intrinsic novelty have appeared in both fields – e.g. Schmidhuber1991Possibility’s theory of curiosity-driven learning (rewarding novelty or surprise), Pathak2017Curiositydriven’s curiosity modules in reinforcement learning, or Secretan2008Picbreeder’s interactive evolution of images – but such metrics only capture narrow aspects of what humans find creative. Purely quantitative measures of complexity or interestingness often fail to match the nuanced human sense of what is truly creative. The challenge, then, is not only to generate novelty, but to ensure an interestingness that is sustainable and meaningful. Recent work argues that open-endedness might be essential for advanced AI: Hughes2024Position posit that any artificial super-intelligence must continually self-improve and discover novelty to surpass human abilities. In practical terms, this means moving beyond training on static human datasets. As silver2025welcome observe, high-quality human data is finite and nearly exhausted for current AI systems – progress from human examples is slowing, necessitating a new paradigm. Open-ended systems address this by enabling agents to generate their own experience and thereby their own ever-growing training data. In summary, open-endedness and interestingness form a guiding vision for next-generation AI: systems that, like living or creative processes, can grow autonomously, continually yielding new, interesting outcomes without end.
@@ Results
Every expert we interviewed underscored the need to suspend judgment during improvisation in order to enter a flow state. P4 put it succinctly: “When you’re improvising, you can’t be judging every move as ‘good’ or ‘bad’ – if you do, you hesitate and the flow is gone. You have to treat everything as an offer and run with it.” This ethos of radical acceptance is taught in improv theater as the “Yes, and” rule – you accept whatever your partner or the situation provides, and build upon it. In dance and music, it manifests as not second-guessing oneself. P6 noted that even if she makes an awkward move in contact improv, she immediately “recycles” it into the next movement rather than stopping in embarrassment. This creates an unbroken continuity: the improvisation may twist in unexpected directions, but it doesn’t halt. The concept of flow (Nakamura2014Concept) – a state of absorbed focus where action and awareness merge – came up repeatedly. Improvisers described flow experiences where “time dilates and self-consciousness disappears.” Neuroscience corroborates this: during improvisation, brain regions associated with self-monitoring and evaluation deactivate, supporting a state of free expression (Limb2008Neural). To achieve this, improvisers deliberately quiet their inner critic. P5 said, “I tell my students to improvise as if nobody’s watching, even if on stage – that freedom is where the real ideas come.” Interestingly, this non-judgment extends to how improvisers treat others in group improv: there’s a norm of not criticizing or blocking another’s contribution, but rather finding a way to make it work. This creates a safe space where creativity can flourish without fear.
Implications
@@
Great improvisations often feel like they are walking a tightrope between order and chaos. Our interviewees described it as riding the edge: “If it’s too predictable, it gets boring; if it’s too random, it falls apart,” said P4. The interestingness sweet spot lies in between – what complexity science calls the edge of chaos, where a system has enough stability to maintain coherence but enough instability to generate novelty. P2 (musician) gave a musical example: in jazz, improvisers sometimes go “outside” the chord changes (deliberately dissonant or offbeat) and then resolve back “inside,” creating tension and release. “It’s like we stretch the rubber band of the music but don’t let it snap,” he explained. Similarly, P1 (dancer) mentioned playing with the boundaries of control: moving just beyond her comfort zone (faster, or with a new quality) so that she has to adapt creatively, but not so far that she loses all technique. Improvisers also often impose light constraints or structures on themselves to shape the chaos. P3 (live coder) might decide to restrict to a certain scale or algorithm during a segment, using that as a launchpad for creativity. These self-imposed rules create a framework that prevents total randomness while still demanding invention within new bounds (a concept analogous to exploring adjacent possible in creative theory).
P4 (CI dancer) reframed two guiding principles in CI as self-sustain and reversibility. For self-sustain, he emphasized being “self-sustaining—like wu wei in Daoist philosophy. The structure comes from your body—its natural form—not imposed by external norms or choreography. It’s your responsibility to hold, protect, and sustain your own body, not rely on others.” For reversibility, he proposed that “all your actions should be reversible to prevent entanglement with others; this provides the maximum possibility for the next move and keeps it flowing. Reversibility gives you the maximal possibility for evolution.”
Implications
@@ Conclusion
Improvisation in human artistry offers rich insights for crafting the next generation of interactive, experiential AI systems. These human-centric qualities can inform concrete design principles for AI. In essence, we envision Experiential AI improvisers that are not just content generators but partners—with a degree of presence and creativity that makes interacting with them feel alive, interesting, and continuously meaningful. By grounding future AI improvisers in the hard-won wisdom of ancient improvisational practices and the spirit of life’s improvisational nature, we can move closer to AI that doesn’t merely output on command, but improvises—co-creating endless interesting possibilities hand-in-hand with humans.